OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 5 cos 3θ

OpenStudy (jdoe0001):

lol

OpenStudy (loser66):

read it http://www.most.gov.mm/techuni/media/EM_02011_chap5b.pdf

OpenStudy (jdoe0001):

testing for y-symmetry let's make r = 5 cos 3θ and set \(\theta = \pi - \theta\) so \(\bf r = 5 cos 3\theta\\ \theta = (\pi - \theta)\\ r = 5cos(3(\pi - \theta)) \implies r = 5cos(3\pi- 3\theta) \)

OpenStudy (jdoe0001):

our resultant equation, DOESN'T RESEMBLE the original equation, therefore, no y-symmetry now let's check for x-symmery by setting \(\theta = - \theta\)

OpenStudy (jdoe0001):

hold on, lemme check really quick and expand the cosine above

OpenStudy (jdoe0001):

$$\bf r = 5 cos 3\theta\\ \theta = (\pi - \theta)\\ r = 5cos(3(\pi - \theta)) \implies r = 5cos(3\pi- 3\theta)\\ cos(3\pi- 3\theta) = cos(3\pi)cos(3\theta)+sin(3\pi)sin(3\theta)\\ cos(3\pi- 3\theta) =-cos(3\theta)\\ \text{thus then}\\ r = 5cos(3\pi- 3\theta) = -5cos(3\theta) $$ so our resultant equation, DOESN'T RESEMBLE the original equation, therefore, no y-symmetry

OpenStudy (jdoe0001):

now let's test for x-symmetry by setting \(\theta = -\theta\)

OpenStudy (jdoe0001):

$$\bf r = 5 cos 3\theta\\ \theta = ( - \theta)\\ r = 5cos(3(- \theta)) \implies r = 5cos(-3\theta)\\ \color{blue}{cos(-x) = cos(x)} \text{ thus then}\\ r = 5cos(-3\theta) = r = 5 cos 3\theta $$ so our resultant equation DOES RESEMBLE the original one, thus there's an x-symmetry

OpenStudy (jdoe0001):

so now let's test for the origin by setting "r" to -r so $$\bf r = 5 cos 3\theta\\ r = -r\\ -r = 5 cos 3\theta \implies r = -5 cos 3\theta $$ so our resultant DOESN'T RESEMBLE the original function, thus no dice on the origin symmetry

OpenStudy (anonymous):

thank you soooooo much :)

OpenStudy (jdoe0001):

yw

OpenStudy (jhannybean):

I'm not understanding why \(\theta = \pi - \theta\) ...

OpenStudy (jdoe0001):

to check for symmetry along the y-axis, or the so-called \(\bf\large \frac{\pi}{2}\) line you set the \(\theta = \pi - \theta\) and then factor and simplify to see if you get the original equation

OpenStudy (jdoe0001):

for the test one can use either \(\bf \theta = \pi - \theta\) or you set "r" and \(\theta\) both to \(\bf -r\ and \ -\theta\)

OpenStudy (jdoe0001):

 and \(\bf \large \pi-a\) on the 2nd quadrant, will show symmetry over the y-axis